OpenStudy (playgames1122):

51 participants went on the 7th grade field trip to the natural history museum. 2/3 of them rode on a large bus. the rest went in a van. How many were in each vehicle

OpenStudy (texaschic101):

2/3 of 51....of means multiply 2/3 * 51 = 102/3 = 34...so 34 rode the bus. 51 - 34 = 17 rode the van

OpenStudy (playgames1122):

Thanks

OpenStudy (texaschic101):

yw